Write the sum using summation notation, assuming the suggested pattern continues.

64 + 81 + 100 + 121 + ... + n2 + ...

summation of n squared from n equals eight to infinity

summation of n minus one squared from n equals eight to infinity

summation of n squared from n equals nine to infinity

summation of n plus one squared from n equals eight to infinity

help me and show work and explanation please
1
Expert's answer
2014-05-19T04:17:12-0400
We have the sum
64 + 81 + 100 + 121 + ... + n^2 + ...=8^2+9^2+10^2+11^2+...+n^2+... = summation of n plus one squared from n equals eight to infinity
The first term in this series is 8 that is why we take n from 8.